OpenStudy (anonymous):

find each function, find f(-2), f(-1/2), f(3) and f(7) f(x)=5x+2

OpenStudy (anonymous):

you just plug in the numbers for x and solve. For example: f(-2) = 5(-2) + 2 = -10 + 2 = -8

OpenStudy (anonymous):

I dont get how to do this other one. f(x)=-1/3x+1

OpenStudy (anonymous):

f(3)=5(3)+2=17 f(7)=5(7)+2=37

OpenStudy (anonymous):

f(-1/2)=5(-1/2)+2=-3

OpenStudy (anonymous):

Just substitute the numbers in for x and do the math. -1/3(PutYourNumberHere) + 1

OpenStudy (anonymous):

noras is wrong, f(-1/2) = -1/2.